• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armando Acha

    Log entry time 00:05:45 on October31,2005

    Entry number 158559

    keyword=Shift summary

    Swing Shift:

    Crew:

    SL/TO: Armando Acha
    TO/3rd. Salvatore Frullani

    16:00: Run 6007 half its way. Things running smoothly.

    16:25: Run 6007 done. 30.9K pairs.

    16:46: Called MCC to move BPMB x from -1.217 to -1.400.

    17:21: Run 6008 done. 44.1K pairs.
    18:18: Run 6009 done. 31.9K pairs.

    19:19: Left Arm Trip.IOC Rebooted. We keep running with right arm.

    19:22: Run 6010 done. 33.8K pairs.
    19:25: Run 6011 stopped. 4K pairs.

    20:00: Left arm back. Q2 and Q3 were cycled.

    20:00: Run 6012 stopped. 16.4k pairs. Left arm back now for next run.
    * Pairs obtained from runs 6011 and 6012 must be divided by 2 (right arm only).

    20:58: Run 6013 done. 23.1K pairs.

    21:24: Run 6014 stopped to flip the IHWP (from IN toOUT). 6.5K pairs.
    Compton Run 12581 stopped.

    21:34: IHWP is OUT and Slug number is 17.

    21:44: Run 6015 stopped. MCC PZT test. Compton run 12583 stopped too.
    Run 6015: 3.6K pairs.

    22:02: PZT test done. Raster sie wil be checked at 5uA.

    22:08: Raster size fine. Back to production!! :-)

    22:35: Compton Run 12584 stopped. ROC 5 failed.

    22:41: Run 6016: Full BPM saturation observed after PZT test.
    Stopped at 23:05. Almost no pairs. Saturation problem fixed.
    new run will be started. Thank you Brad.

    00:00 Run 6017 done. 44.6K pairs.

    Some problems prevented us from having a very producitve shift. We lost left arm for a while and unstable beam. Some saturation was observed also in run 6016 but finally everything was fixed and things were left running smoothly.